# Mitigating Label Noise using Prompt-Based Hyperbolic Meta-Learning in Open-Set Domain Generalization

### v1

@@ 5.1 Implementation Details
TCL [26] 22.70 25.47 16.55 22.26 23.14 14.42 19.75 14.23 6.53 28.52 10.92 11.54 23.31 18.44 12.26
NPN [55] 15.27 15.29 7.9 16.07 10.88 11.29 19.96 17.19 8.06 20.48 1.12 18.43 17.95 11.12 11.42
BadLabel [75] 13.33 20.96 13.32 24.77 32.36 21.98 23.98 32.76 21.91 21.60 2.36 21.31 20.92 22.11 19.63
DISC [43] 22.37 1.92 11.07 24.58 19.66 12.81 24.08 13.40 14.20 20.03 4.56 12.61 22.77 9.89 12.67
LSL [29] 10.02 3.64 5.89 26.27 10.83 17.29 31.87 27.63 19.75 25.39 8.09 7.83 23.39 12.55 12.69
PLM [78] 29.32 5.93 1.34 26.77 28.58 18.39 25.27 25.69 16.89 24.22 3.33 2.25 26.4 15.88 9.72
ARPL [3] 25.85 20.35 21.54 20.14 18.60 13.63 24.24 23.62 14.23 16.26 3.36 9.32 21.62 16.48 14.68
ODGNet [4] 23.18 19.01 10.04 22.64 21.69 10.84 19.80 19.01 10.58 18.07 0.60 1.60 20.92 15.08 8.27
MLDG [56] 29.08 29.84 22.82 30.08 29.62 20.16 24.24 22.37 15.85 19.66 2.34 4.11 25.77 21.04 15.74
SWAD [6] 19.42 15.01 15.83 23.83 22.54 14.46 28.14 26.18 16.29 23.59 14.73 9.01 23.75 19.62 13.90
MixStyle [83] 22.70 25.94 18.23 17.32 8.52 15.69 34.97 24.21 16.86 20.99 1.21 10.62 24.00 14.97 15.35
MEDIC-cls [66] 35.70 23.85 14.91 28.46 30.78 20.50 30.22 26.39 16.88 21.51 11.08 13.13 28.97 23.03 16.36
MEDIC-bcls [66] 35.70 23.85 18.53 28.46 27.55 17.87 30.22 13.99 11.18 21.54 8.32 4.81 28.98 18.43 13.10
@@
Table 8: Results (%) of PACS on ViT-Base [16]. The open-set ratio is $6{:}1$ and asymmetric label noise with ratio $50\%$ is selected.
@@
TCL [26] 51.21 56.51 48.47 46.65 51.35 40.80 31.92 30.94 22.61 25.87 20.94 12.04 38.91 39.94 30.98
NPN [55] 27.63 8.86 6.32 32.40 19.00 15.72 21.66 16.87 14.25 20.48 26.15 16.50 25.54 17.72 13.20
BadLabel [75] 46.12 58.96 45.77 35.33 34.30 34.83 26.92 39.52 26.18 20.70 31.32 20.66 32.27 41.03 31.86
DISC [43] 47.17 17.08 9.77 24.64 19.96 10.75 20.22 12.84 9.22 23.93 4.31 15.10 28.99 13.55 11.21
LSL [29] 49.52 5.63 15.99 37.52 34.68 26.68 28.16 27.60 18.13 28.16 27.60 18.13 35.84 23.88 19.73
PLM [78] 20.51 23.01 18.15 20.17 13.41 8.08 37.85 33.88 26.43 20.59 24.16 17.26 24.78 23.62 17.48
ARPL [3] 50.32 53.47 44.10 44.03 42.48 33.24 43.15 36.75 30.20 17.72 13.62 12.77 38.81 36.58 30.08
ODGNet [4] 51.13 55.83 48.95 49.16 45.58 37.21 40.69 37.35 28.85 21.76 17.55 12.68 40.69 39.08 31.92
MLDG [56] 52.99 57.10 51.53 45.90 56.06 43.93 46.11 43.18 39.17 32.98 29.49 28.18 44.50 46.46 40.70
SWAD [6] 50.32 54.99 45.45 44.84 54.03 42.11 39.92 45.57 35.45 25.42 18.78 12.07 40.13 43.34 33.77
MixStyle [83] 53.72 53.47 52.97 46.15 48.22 39.13 44.87 44.84 36.37 29.50 12.81 19.03 43.56 39.84 36.88
MEDIC-cls [66] 52.10 59.31 48.42 46.47 56.57 44.42 35.69 29.04 21.31 35.42 28.01 27.79 42.42 43.23 35.49
MEDIC-bcls [66] 52.10 49.72 42.70 46.47 55.52 43.88 35.69 30.26 20.16 35.42 24.54 26.48 42.42 40.01 33.31
@@
[…]
@@
NPN [55] 82.28 29.42 70.68 32.56 28.22 22.74 21.78 22.54 14.41 17.33 11.95 9.89 38.49 23.03 29.43
BadLabel [75] 63.25 52.49 61.18 30.00 34.82 27.42 50.17 51.19 46.06 41.82 16.75 41.82 46.31 38.81 44.12
ODGNet [4] 90.33 50.84 71.38 59.28 26.10 43.83 70.11 53.83 49.54 55.81 32.71 39.54 68.88 40.87 51.07
MLDG [7] 90.67 27.60 80.46 57.89 48.38 42.77 60.33 39.95 41.30 53.75 6.08 35.20 65.66 30.50 49.93
MEDIC-cls [66] 22.08 8.71 5.89 21.33 20.09 9.94 23.24 9.83 13.28 13.72 10.29 6.63 20.09 12.23 8.94
MEDIC-bcls [66] 22.08 12.31 5.57 21.33 16.19 10.11 23.24 12.64 5.41 13.72 13.47 5.94 20.09 13.65 6.76
@@
Table 10: Results (%) of DigitsDG on ConvNet [84], where symmetric label noise with ratio $50\%$ is selected.
@@
NPN [55] 68.11 25.80 49.14 28.31 28.29 18.65 45.78 32.17 31.15 28.42 1.38 16.37 42.66 21.91 28.83
BadLabel [75] 63.31 35.12 57.47 42.28 42.69 38.44 21.36 25.07 18.30 19.83 4.88 19.28 36.70 26.94 33.37
ODGNet [4] 71.25 22.26 49.81 59.22 36.88 44.86 61.39 43.65 42.41 51.75 19.70 35.38 60.90 30.62 43.12
MLDG [56] 62.72 54.21 50.52 48.94 41.54 35.35 43.53 33.15 29.09 24.61 6.48 14.29 44.95 33.85 32.31
MEDIC-cls [66] 22.39 0.17 2.81 25.89 22.88 12.16 12.78 12.10 4.73 7.61 5.70 2.35 17.17 10.21 5.51
MEDIC-bcls [66] 22.39 12.08 3.26 25.89 20.73 11.88 12.78 11.26 3.97 7.61 7.13 2.02 17.17 12.80 5.28
@@
Table 11: Results (%) of DigitsDG on ConvNet [84], where symmetric label noise with ratio $80\%$ is selected..
@@
NPN [55] 16.67 0.01 9.65 18.61 13.38 10.15 17.78 18.12 9.29 16.67 0.55 8.21 17.43 8.02 9.33
BadLabel [75] 18.58 5.84 7.39 17.39 15.13 8.99 16.67 0.41 8.36 17.11 0.22 8.44 17.44 5.40 8.30
ODGNet [4] 16.19 1.51 10.37 17.28 11.80 10.64 18.47 17.04 8.49 16.72 14.89 8.59 17.17 11.31 9.52
MLDG [56] 16.06 6.70 9.69 18.58 3.26 9.27 16.94 6.81 8.20 17.33 6.88 8.84 17.23 5.91 9.00
MEDIC-cls [66] 21.17 3.51 11.37 18.75 16.52 7.77 15.81 4.48 7.54 17.11 4.86 8.64 18.21 7.34 8.83
MEDIC-bcls [66] 21.17 7.21 8.32 16.83 13.25 8.81 15.81 4.44 7.70 17.11 4.96 8.38 17.73 7.47 8.30
@@
Table 12: Results (%) of DigitsDG on ConvNet [84], where asymmetric label noise with ratio $50\%$ is selected.
@@
NPN [55] 71.08 24.46 61.45 54.58 41.21 38.79 55.92 42.95 35.03 51.97 11.28 34.23 58.39 29.98 42.38
BadLabel [75] 53.00 38.77 43.34 33.94 27.79 30.06 37.64 44.59 33.47 27.61 20.90 26.18 38.05 33.01 33.26
ODGNet [4] 53.67 48.23 41.31 39.36 27.19 28.08 51.17 24.92 34.91 39.64 5.51 25.32 45.96 26.46 32.41
MLDG [56] 68.17 23.34 55.68 56.47 40.89 40.97 56.31 41.94 38.02 47.81 11.35 31.60 57.19 29.38 41.57
MEDIC-cls [66] 19.86 14.41 11.22 19.75 15.93 8.03 9.83 8.28 2.88 16.31 5.78 7.36 16.44 11.10 7.37
MEDIC-bcls [66] 19.86 21.31 11.22 19.75 14.66 6.73 9.83 8.82 2.58 16.31 13.40 7.39 16.44 14.55 6.98
@@ 5.2 Analysis of the Benchmarks
Analysis of benchmark on PACS: We construct our benchmark according to the aforementioned label noise setting in Section 3.2 on both of the PACS and DigitsDG datasets, where we deploy ResNet18 [24] (Table 1 to Table 4) and ViT-Base [16] (Table 5 to Table 8) on the PACS dataset and ConvNet [84] on the DigitsDG dataset (Table 9 to Table 12).
@@
First, we observe that with the increase of the label noise ratio, there are degradations of the OSDG performance on all the leveraged baselines, according to the Table 1, Table 2, and Table 3, where symmetric label noise with ratio $20\%$, $50\%$, and $80\%$ are used on PACS and ResNet18 [24]. Among all the approaches derived from the label noise learning field, BadLabel [75], NPN [55], and DISC [43] show promising performances when tackling OSDG-NL task, where BadLabel [75] delivers $47.61\%$, $38.78\%$, $21.02\%$, and $27.26\%$ of closed-set accuracy, $46.55\%$, $43.76\%$, $16.77\%$, and $31.82\%$ in terms of H-score, and $42.22\%$, $32.70\%$, $16.42\%$, and $26.54\%$ in terms of OSCR on PACS with symmetric label noise ratio $20\%$ (Table 1), $50\%$ (Table 2), $80\%$ (Table 3) and asymmetric label noise ratio $50\%$ (Table 4) on ResNet18 [24] backbone, respectively.
@@
Among all the approaches from the OSDG field, GAN-based method, i.e., ODGNet [4], and meta-learning-based methods, i.e., MLDG [56] and MEDIC [66], show good performance when dealing with various adopted label noises, where MEDIC [66] delivers $52.66\%$, $50.27\%$, $26.40\%$, and $37.79\%$ of closed-set accuracy, $49.25\%$, $40.24\%$, $16.25\%$, and $31.73\%$ in terms of H-score, and $42.76\%$, $34.62\%$, $11.17\%$, and $23.95\%$ in terms of OSCR on PACS with symmetric label noise ratio $20\%$ (Table 1), $50\%$ (Table 2), $80\%$ (Table 3) and asymmetric label noise ratio $50\%$ (Table 4) on the ResNet18 backbone using binary classification head for each category (bcls), respectively. The label noise learning approach, BadLabel [75], outperforms OSDG approaches in the scenario where a large label noise ratio is adopted. Similar observations can also be found in the experiments conducted with ViT-Base [16] backbone shown in Table 5 to Table 8.
@@
When we compare Table 2 and Table 4, we find out that asymmetric label noise is much more challenging compared with the symmetric label noise under the same label noise ratio $50\%$. Among all the leveraged baselines using ResNet18 [24] as the backbone, DISC [43] shows $30.47\%$ in terms of OSCR under asymmetric label noise but still yields $0.94\%$ performance degradation when compared with the symmetric scenario. MLDG [7] achieves better performance compared with the other leveraged baselines on ViT-Base [16] backbone when facing with asymmetric label noise with ratio $50\%$, where $40.70\%$ in terms of OSCR is delivered, as shown in Table 8. The aforementioned observation highlights the challenge posed by label noise in the OSDG task, a problem that existing label noise learning methods and OSDG approaches have yet to effectively address.
@@
The proposed approach, HyProMeta, surpasses all baseline methods by employing label noise-agnostic meta-learning and new-category-aware prompt learning. Specifically, categorical prototypes derived from the hyperbolic space guide data partitioning during the meta-task assignment, while learnable prompt enhances mixed-category samples, thereby improving the generalizability of the resulting model. HyProMeta achieves superior performance, outperforming the strongest baselines by significant margins across various backbones and label noise ratios on PACS. For instance, under a symmetric label noise ratio of $20\%$, HyProMeta surpasses NPN [55] by $4.32\%$, $8.69\%$, and $5.23\%$ in terms of closed-set accuracy, H-score, and OSCR, respectively, when using ResNet18 [24] as the feature learning backbone, as shown in Table 1.
@@
When adopting a more powerful backbone, i.e., ViT-Base [16], both the baseline methods and our proposed approach achieve further performance improvements. Specifically, HyProMeta achieves closed-set accuracy of $59.65\%$, $58.68\%$, $37.06\%$, and $49.99\%$; H-score of $60.06\%$, $59.33\%$, $29.09\%$, and $48.47\%$; and OSCR of $54.97\%$, $52.91\%$, $25.26\%$, and $43.44\%$ on PACS using ViT-Base [16] as the backbone, significantly outperforming existing approaches, as presented in Table 5 to Table 8.
@@
Analysis of benchmark on DigitsDG: To further assess the cross-dataset generalizability of our proposed approach and the baselines for the OSDG-NL task, we selected several baselines that demonstrated promising performance on the PACS dataset to construct a benchmark on another OSDG dataset, i.e., DigitsDG. The results, presented in Table 9 to Table 12, consider symmetric label noise with ratios $20\%$, $50\%$, and $80\%$, as well as asymmetric label noise with a ratio of $50\%$.
We first observe that the MEDIC approach [66] fails to perform on DigitsDG when using ConvNet [84] as the backbone. This demonstrates that MEDIC relies heavily on high-quality a priori information from the source domains to ensure its effectiveness, particularly on small datasets and when employing lightweight backbones. Among all the baselines evaluated on the DigitsDG dataset, ODGNet [4] achieves the best performance under symmetric label noise, while NPN [55] demonstrates superior performance under asymmetric label noise.
Table 13: Module ablation on PACS using ResNet18 [24] under symmetric label noise ratio $50\%$.
@@
Table 14: Ablation study of NCA-Prompt on symmetric label noise $50\%$ on PACS dataset using ResNet18 [24].
@@
Table 15: Ablation study of HYB-Meta on symmetric label noise $50\%$ on PACS dataset using ResNet18 [24].

### v2

@@ 5.2 Evaluation Metrics
TCL [huang2023twin] 22.70 25.47 16.55 22.26 23.14 14.42 19.75 14.23 6.53 28.52 10.92 11.54 23.31 18.44 12.26
NPN [sheng2024adaptive] 15.27 15.29 7.9 16.07 10.88 11.29 19.96 17.19 8.06 20.48 1.12 18.43 17.95 11.12 11.42
BadLabel [zhang2024badlabel] 13.33 20.96 13.32 24.77 32.36 21.98 23.98 32.76 21.91 21.60 2.36 21.31 20.92 22.11 19.63
DISC [li2023disc] 22.37 1.92 11.07 24.58 19.66 12.81 24.08 13.40 14.20 20.03 4.56 12.61 22.77 9.89 12.67
LSL [kim2024learning] 10.02 3.64 5.89 26.27 10.83 17.29 31.87 27.63 19.75 25.39 8.09 7.83 23.39 12.55 12.69
PLM [zhao2024estimating] 29.32 5.93 1.34 26.77 28.58 18.39 25.27 25.69 16.89 24.22 3.33 2.25 26.4 15.88 9.72
ARPL [bendale2016towards] 25.85 20.35 21.54 20.14 18.60 13.63 24.24 23.62 14.23 16.26 3.36 9.32 21.62 16.48 14.68
ODGNet [bose2023beyond] 23.18 19.01 10.04 22.64 21.69 10.84 19.80 19.01 10.58 18.07 0.60 1.60 20.92 15.08 8.27
MLDG [shu2019meta] 29.08 29.84 22.82 30.08 29.62 20.16 24.24 22.37 15.85 19.66 2.34 4.11 25.77 21.04 15.74
SWAD [cha2021swad] 19.42 15.01 15.83 23.83 22.54 14.46 28.14 26.18 16.29 23.59 14.73 9.01 23.75 19.62 13.90
MixStyle [zhou2020domain] 22.70 25.94 18.23 17.32 8.52 15.69 34.97 24.21 16.86 20.99 1.21 10.62 24.00 14.97 15.35
MEDIC-cls [wang2023generalizable] 35.70 23.85 14.91 28.46 30.78 20.50 30.22 26.39 16.88 21.51 11.08 13.13 28.97 23.03 16.36
MEDIC-bcls [wang2023generalizable] 35.70 23.85 18.53 28.46 27.55 17.87 30.22 13.99 11.18 21.54 8.32 4.81 28.98 18.43 13.10
EBiL-HaDS-cls [peng2024advancing] 34.89 3.49 17.80 32.02 20.65 22.32 30.84 6.28 17.64 25.71 26.60 17.09 30.87 14.26 18.71
EBiL-HaDS-bcls [peng2024advancing] 34.89 20.63 16.79 32.02 26.79 17.85 30.84 24.30 16.61 25.71 29.67 16.33 30.87 25.35 16.90
@@
Table 8: Results (%) of PACS on ViT-Base [dosovitskiy2021an]. The open-set ratio is 6:16{:}1 and asymmetric label noise with a ratio of $50\%$ is selected.
@@
TCL [huang2023twin] 51.21 56.51 48.47 46.65 51.35 40.80 31.92 30.94 22.61 25.87 20.94 12.04 38.91 39.94 30.98
NPN [sheng2024adaptive] 27.63 8.86 6.32 32.40 19.00 15.72 21.66 16.87 14.25 20.48 26.15 16.50 25.54 17.72 13.20
BadLabel [zhang2024badlabel] 46.12 58.96 45.77 35.33 34.30 34.83 26.92 39.52 26.18 20.70 31.32 20.66 32.27 41.03 31.86
DISC [li2023disc] 47.17 17.08 9.77 24.64 19.96 10.75 20.22 12.84 9.22 23.93 4.31 15.10 28.99 13.55 11.21
LSL [kim2024learning] 49.52 5.63 15.99 37.52 34.68 26.68 28.16 27.60 18.13 28.16 27.60 18.13 35.84 23.88 19.73
PLM [zhao2024estimating] 20.51 23.01 18.15 20.17 13.41 8.08 37.85 33.88 26.43 20.59 24.16 17.26 24.78 23.62 17.48
ARPL [bendale2016towards] 50.32 53.47 44.10 44.03 42.48 33.24 43.15 36.75 30.20 17.72 13.62 12.77 38.81 36.58 30.08
ODGNet [bose2023beyond] 51.13 55.83 48.95 49.16 45.58 37.21 40.69 37.35 28.85 21.76 17.55 12.68 40.69 39.08 31.92
MLDG [shu2019meta] 52.99 57.10 51.53 45.90 56.06 43.93 46.11 43.18 39.17 32.98 29.49 28.18 44.50 46.46 40.70
SWAD [cha2021swad] 50.32 54.99 45.45 44.84 54.03 42.11 39.92 45.57 35.45 25.42 18.78 12.07 40.13 43.34 33.77
MixStyle [zhou2020domain] 53.72 53.47 52.97 46.15 48.22 39.13 44.87 44.84 36.37 29.50 12.81 19.03 43.56 39.84 36.88
MEDIC-cls [wang2023generalizable] 52.10 59.31 48.42 46.47 56.57 44.42 35.69 29.04 21.31 35.42 28.01 27.79 42.42 43.23 35.49
MEDIC-bcls [wang2023generalizable] 52.10 49.72 42.70 46.47 55.52 43.88 35.69 30.26 20.16 35.42 24.54 26.48 42.42 40.01 33.31
EBiL-HaDS-cls [peng2024advancing] 54.60 52.92 44.98 46.97 56.52 44.29 34.66 25.52 22.48 25.71 26.60 17.09 40.49 40.39 32.21
EBiL-HaDS-bcls [peng2024advancing] 54.60 39.01 29.61 46.97 54.34 42.74 34.66 31.86 22.47 25.71 29.67 16.33 40.49 38.72 27.79
@@
Table 9: Results ($\%$) of DigitsDG on ConvNet [zhou2021domain], where symmetric label noise with a ratio of $20\%$ is selected.
@@
NPN [sheng2024adaptive] 82.28 29.42 70.68 32.56 28.22 22.74 21.78 22.54 14.41 17.33 11.95 9.89 38.49 23.03 29.43
BadLabel [zhang2024badlabel] 63.25 52.49 61.18 30.00 34.82 27.42 50.17 51.19 46.06 41.82 16.75 41.82 46.31 38.81 44.12
ODGNet [bose2023beyond] 90.33 50.84 71.38 59.28 26.10 43.83 70.11 53.83 49.54 55.81 32.71 39.54 68.88 40.87 51.07
MLDG [chen2021adversarial] 90.67 27.60 80.46 57.89 48.38 42.77 60.33 39.95 41.30 53.75 6.08 35.20 65.66 30.50 49.93
MEDIC-cls [wang2023generalizable] 22.08 8.71 5.89 21.33 20.09 9.94 23.24 9.83 13.28 13.72 10.29 6.63 20.09 12.23 8.94
MEDIC-bcls [wang2023generalizable] 22.08 12.31 5.57 21.33 16.19 10.11 23.24 12.64 5.41 13.72 13.47 5.94 20.09 13.65 6.76
EBiL-HaDS-cls [peng2024advancing] 88.28 48.19 78.11 42.86 33.81 29.72 72.36 54.16 51.49 52.33 39.08 38.24 63.96 43.81 49.39
EBiL-HaDS-bcls [peng2024advancing] 88.28 59.51 61.82 42.86 34.27 30.03 72.36 53.55 49.29 52.33 35.23 40.31 63.96 45.64 45.36
@@
Table 10: Results (%) of DigitsDG on ConvNet [zhou2021domain], where symmetric label noise with a ratio of $50\%$ is selected.
@@
NPN [sheng2024adaptive] 68.11 25.80 49.14 28.31 28.29 18.65 45.78 32.17 31.15 28.42 1.38 16.37 42.66 21.91 28.83
BadLabel [zhang2024badlabel] 63.31 35.12 57.47 42.28 42.69 38.44 21.36 25.07 18.30 19.83 4.88 19.28 36.70 26.94 33.37
ODGNet [bose2023beyond] 71.25 22.26 49.81 59.22 36.88 44.86 61.39 43.65 42.41 51.75 19.70 35.38 60.90 30.62 43.12
MLDG [shu2019meta] 62.72 54.21 50.52 48.94 41.54 35.35 43.53 33.15 29.09 24.61 6.48 14.29 44.95 33.85 32.31
MEDIC-cls [wang2023generalizable] 22.39 0.17 2.81 25.89 22.88 12.16 12.78 12.10 4.73 7.61 5.70 2.35 17.17 10.21 5.51
MEDIC-bcls [wang2023generalizable] 22.39 12.08 3.26 25.89 20.73 11.88 12.78 11.26 3.97 7.61 7.13 2.02 17.17 12.80 5.28
EBiL-HaDS-cls [peng2024advancing] 77.19 31.34 62.64 42.78 34.74 32.69 51.92 39.84 31.75 43.81 23.45 30.23 53.93 32.34 39.33
EBiL-HaDS-bcls [peng2024advancing] 77.19 47.43 47.86 42.78 30.28 22.24 51.92 18.79 22.75 43.81 30.42 20.81 53.93 31.73 28.42
@@
Table 11: Results (%) of DigitsDG on ConvNet [zhou2021domain], where symmetric label noise with a ratio of $80\%$ is selected..
@@
NPN [sheng2024adaptive] 16.67 0.01 9.65 18.61 13.38 10.15 17.78 18.12 9.29 16.67 0.55 8.21 17.43 8.02 9.33
BadLabel [zhang2024badlabel] 18.58 5.84 7.39 17.39 15.13 8.99 16.67 0.41 8.36 17.11 0.22 8.44 17.44 5.40 8.30
ODGNet [bose2023beyond] 16.19 1.51 10.37 17.28 11.80 10.64 18.47 17.04 8.49 16.72 14.89 8.59 17.17 11.31 9.52
MLDG [shu2019meta] 16.06 6.70 9.69 18.58 3.26 9.27 16.94 6.81 8.20 17.33 6.88 8.84 17.23 5.91 9.00
MEDIC-cls [wang2023generalizable] 21.17 3.51 11.37 18.75 16.52 7.77 15.81 4.48 7.54 17.11 4.86 8.64 18.21 7.34 8.83
MEDIC-bcls [wang2023generalizable] 21.17 7.21 8.32 16.83 13.25 8.81 15.81 4.44 7.70 17.11 4.96 8.38 17.73 7.47 8.30
EBiL-HaDS-cls [peng2024advancing] 12.72 7.25 5.36 16.14 12.10 8.46 16.92 12.83 8.17 16.44 5.27 7.79 15.56 9.36 7.45
EBiL-HaDS-bcls [peng2024advancing] 12.72 7.15 5.62 16.14 8.64 7.43 16.92 15.29 8.34 16.44 7.00 7.83 15.56 9.52 7.31
@@
Table 12: Results (%) of DigitsDG on ConvNet [zhou2021domain], where asymmetric label noise with a ratio of $50\%$ is selected.
@@
NPN [sheng2024adaptive] 71.08 24.46 61.45 54.58 41.21 38.79 55.92 42.95 35.03 51.97 11.28 34.23 58.39 29.98 42.38
BadLabel [zhang2024badlabel] 53.00 38.77 43.34 33.94 27.79 30.06 37.64 44.59 33.47 27.61 20.90 26.18 38.05 33.01 33.26
ODGNet [bose2023beyond] 53.67 48.23 41.31 39.36 27.19 28.08 51.17 24.92 34.91 39.64 5.51 25.32 45.96 26.46 32.41
MLDG [shu2019meta] 68.17 23.34 55.68 56.47 40.89 40.97 56.31 41.94 38.02 47.81 11.35 31.60 57.19 29.38 41.57
MEDIC-cls [wang2023generalizable] 19.86 14.41 11.22 19.75 15.93 8.03 9.83 8.28 2.88 16.31 5.78 7.36 16.44 11.10 7.37
MEDIC-bcls [wang2023generalizable] 19.86 21.31 11.22 19.75 14.66 6.73 9.83 8.82 2.58 16.31 13.40 7.39 16.44 14.55 6.98
EBiL-HaDS-cls [peng2024advancing] 67.39 36.21 53.76 44.14 35.30 29.61 60.86 46.15 38.81 51.89 25.34 31.35 56.07 35.75 38.38
EBiL-HaDS-bcls [peng2024advancing] 67.39 44.88 50.01 44.14 36.20 37.31 60.86 48.40 41.70 51.89 23.25 36.46 56.07 38.18 41.37
@@ 5.3 Analysis on the Benchmarks
Analysis on benchmark on PACS: We construct our benchmark according to the aforementioned label noise setting in Section 3.2 on both of the PACS and DigitsDG datasets, where we deploy ResNet18 [he2016deep] (Table 1 to Table 4) and ViT-Base [dosovitskiy2021an] (Table 5 to Table 8) on the PACS dataset and ConvNet [zhou2021domain] on the DigitsDG dataset (Table 9 to Table 12).
@@
First, we observe that with the increase of the label noise ratio, there are degradations of the OSDG performance on all the leveraged baselines, according to the Table 1, Table 2, and Table 3, where symmetric label noise with ratios of $20\%$, $50\%$, and $80\%$ are used on PACS and ResNet18 [he2016deep]. Among all the approaches derived from the label noise learning field, BadLabel [zhang2024badlabel], NPN [sheng2024adaptive], and DISC [li2023disc] show promising performances when tackling OSDG-NL task, where BadLabel [zhang2024badlabel] delivers $47.61\%$, $38.78\%$, $21.02\%$, and $27.26\%$ of closed-set accuracy, $46.55\%$, $43.76\%$, $16.77\%$, and $31.82\%$ in terms of H-score, and $42.22\%$, $32.70\%$, $16.42\%$, and $26.54\%$ in terms of OSCR on PACS with symmetric label noise ratio $20\%$ (Table 1), $50\%$ (Table 2), $80\%$ (Table 3) and asymmetric label noise ratio $50\%$ (Table 4) on ResNet18 [he2016deep] backbone, respectively.
@@
Among all the approaches from the OSDG field, GAN-based method, i.e., ODGNet [bose2023beyond], and meta-learning-based methods, i.e., MLDG [shu2019meta] and MEDIC [wang2023generalizable], show good performance when dealing with various adopted label noises, where MEDIC [wang2023generalizable] delivers $52.66\%$, $50.27\%$, $26.40\%$, and $37.79\%$ of closed-set accuracy, $49.25\%$, $40.24\%$, $16.25\%$, and $31.73\%$ in terms of H-score, and $42.76\%$, $34.62\%$, $11.17\%$, and $23.95\%$ in terms of OSCR on PACS with symmetric label noise ratio $20\%$ (Table 1), $50\%$ (Table 2), $80\%$ (Table 3) and asymmetric label noise ratio $50\%$ (Table 4) on the ResNet18 backbone using binary classification head for each category (bcls), respectively. EBiL-HaDS-cls [peng2024advancing] and EBiL-HaDS-bcls [peng2024advancing] perform reasonably well under $20\%$ and $50\%$ symmetric noise, particularly in the Photo domain, but both variants exhibit notable degradation as noise severity increases. Under $80\%$ symmetric noise and $50\%$ asymmetric noise, their average OSCR values fall below $28\%$, indicating limited robustness in high-noise scenarios.
The label noise learning approach, BadLabel [zhang2024badlabel], outperforms OSDG approaches in the scenario where a large label noise ratio is adopted. Similar observations can also be found in the experiments conducted with ViT-Base [dosovitskiy2021an] backbone shown in Table 5 to Table 8.
@@
When we compare Table 2 and Table 4, we find that asymmetric label noise is much more challenging compared with the symmetric label noise under the same label noise ratio $50\%$. Among all the leveraged baselines using ResNet18 [he2016deep] as the backbone, DISC [li2023disc] shows $30.47\%$ in terms of OSCR under asymmetric label noise but still yields $0.94\%$ performance degradation when compared with the symmetric scenario. MLDG [chen2021adversarial] achieves better performance compared with the other leveraged baselines on ViT-Base [dosovitskiy2021an] backbone when facing asymmetric label noise with a ratio of $50\%$, where $40.70\%$ in terms of OSCR is delivered, as shown in Table 8. The aforementioned observation highlights the challenge posed by label noise in the OSDG task, a problem that existing label noise learning methods and OSDG approaches have yet to effectively address.
@@
The proposed approach, HyProMeta, surpasses all baseline methods by employing label noise-agnostic meta-learning and new-category-aware prompt learning. Specifically, category prototypes derived from the hyperbolic space guide data partitioning during the meta-task assignment, while learnable prompt enhances mixed-category samples, thereby improving the generalizability of the resulting model. HyProMeta achieves superior performance, outperforming the strongest baselines by significant margins across various backbones and label noise ratios on PACS. For instance, under a symmetric label noise ratio of $20\%$, HyProMeta surpasses NPN [sheng2024adaptive] by $4.32\%$, $8.69\%$, and $5.23\%$ in terms of closed-set accuracy, H-score, and OSCR, respectively, when using ResNet18 [he2016deep] as the feature learning backbone, as shown in Table 1. HyProMeta consistently outperforms both EBiL-HaDS-cls [peng2024advancing] and EBiL-HaDS-bcls [peng2024advancing] across all noise settings on PACS with ResNet18 [he2016deep], achieving the highest average OSCR under $20\%$, $50\%$, $80\%$ symmetric, and $50\%$ asymmetric noise, with respective margins of up to $5\%$ in terms of OSCR. The performance gap becomes especially prominent as noise severity increases, demonstrating HyProMeta’s superior robustness to label corruption and domain shift.
@@
When adopting a more powerful backbone, i.e., ViT-Base [dosovitskiy2021an], both the baseline methods and our proposed approach achieve further performance improvements. Specifically, HyProMeta achieves closed-set accuracy of $59.65\%$, $58.68\%$, $37.06\%$, and $49.99\%$; H-score of $60.06\%$, $59.33\%$, $29.09\%$, and $48.47\%$; and OSCR of $54.97\%$, $52.91\%$, $25.26\%$, and $43.44\%$ on PACS using ViT-Base [dosovitskiy2021an] as the backbone, significantly outperforming existing approaches, as presented in Table 5 to Table 8.
@@
Analysis on benchmark on DigitsDG: To further assess the cross-dataset generalizability of our proposed approach and the baselines for the OSDG-NL task, we selected several baselines that demonstrated promising performance on the PACS dataset to construct a benchmark on another OSDG dataset, i.e., DigitsDG. The results, presented in Table 9 to Table 12, consider symmetric label noise with ratios of $20\%$, $50\%$, and $80\%$, as well as asymmetric label noise with a ratio of $50\%$.
We first observe that the MEDIC approach [wang2023generalizable] fails to perform well on DigitsDG when using ConvNet [zhou2021domain] as the backbone. This demonstrates that MEDIC relies heavily on high-quality cues from the source domains to ensure its effectiveness, particularly on small datasets and when employing lightweight backbones. Among all the baselines evaluated on the DigitsDG dataset, ODGNet [bose2023beyond] achieves the best performance under symmetric label noise, while NPN [sheng2024adaptive] demonstrates superior performance under asymmetric label noise.
Table 13: Module ablation on PACS using ResNet18 [he2016deep] under symmetric label noise ratio $50\%$.
@@
Table 14: Ablation study of NCA-Prompt on symmetric label noise $50\%$ on PACS dataset using ResNet18 [he2016deep].
@@
Table 15: Ablation study of HYB-Meta on symmetric label noise $50\%$ on PACS dataset using ResNet18 [he2016deep].